OpenStudy (king_lexx35):

ANSWER GETS FAN AND METAL Which expression is equal to sec x° multiplied by cot x°? hypotenuse ÷ adjacent adjacent ÷ hypotenuse opposite ÷ adjacent hypotenuse ÷ opposite

TheSmartOne (thesmartone):

Hint: sec x = 1/cos x cot x = 1/tan x sin x = opp/hyp cos x = adj/hyp tan x = sin(x)/cos(x)

OpenStudy (king_lexx35):

WHATS THE ANSWER PLEASE @TheSmartOne

TheSmartOne (thesmartone):

On OpenStudy, we do not give answers. We guide users to the answer.

TheSmartOne (thesmartone):

I was just about to tell you how you could get the answer, but alrighty.

jhonyy9 (jhonyy9):

so than you know that sec x = 1/cos x and that cot x = 1/tan x = 1/(sinx/cos x) so in this way you get cot x = cos x /sin x than you rewrite it now the sec x * cot x = ? what will get using these above wrote ?

jhonyy9 (jhonyy9):

please collaborate bc. you need getting the roght answer with our help exactly how wrote to you above @TheSmartOne

OpenStudy (king_lexx35):

YOU CANT JUST GIVE THE ANSWER? @jhonyy9

jhonyy9 (jhonyy9):

no bc. the rule of OS. interdict and bc. you are here to learn how you can solve the same exercices not just for get the right answers - sorry

OpenStudy (king_lexx35):

OHH IM SORRY I DID NOT KNOW THAT @jhonyy9 @TheSmartOne

OpenStudy (king_lexx35):

I THOUGHT YALLL JUST DIDNT WANT TO GIVE ME THE ANSWER @TheSmartOne @jhonyy9